# Discovering Coordinated Joint Options via Inter-Agent Relative Dynamics

### v1

@@ 1 Introduction
In single-agent reinforcement learning (RL), the options framework (sutton1999optionsdef) is a widely used mechanism for formulating temporally extended actions. That is, options can act as shortcuts between distant regions of the state space during exploration (Barto2001). Yet, as noted in prior work (Jong2008TheUO), the effectiveness of options is sensitive to many factors, and poorly designed and excessively large sets of options can hinder learning. This makes option discovery, i.e., the automated design of useful options, a challenging problem. Methods based on the eigen-decomposition of the graph Laplacian, such as Eigenoptions (Machado2017), have gained traction due to their task-agnostic discovery of options and exploration guarantees (Jinnai2019opt). However, the reliance on the eigenvectors of the state-transition graph Laplacian leads to an excessive number of options being discovered (twice the state count), a problem that is more pronounced in multi-agent systems because the joint state space grows exponentially with the number of agents. Moreover, current Laplacian eigenvector approximators are most effective in estimating small numbers of eigenvectors (wang2021optiondisc; gomez2024optiondisc), risking not identifying many useful options, particularly those that facilitate exploration at various timescales.
We address option discovery for multi-agent systems through a novel inter-agent relative state abstraction. This new state abstraction compresses the joint state space of a group of agents into a compact latent representation centred around the state of maximal alignment among agents we call the Fermat state. Through this abstraction of the joint state, we drastically reduce the number of discovered options while also focusing the discovery process on inter-agent relational dynamics. We then empirically show how this transformation encourages the emergence of highly coordinated behaviours. Our approach builds on the intuition that in the absence of an explicit objective, synchronisation over state features represents a natural basis for coordination. Returning to the ball game example, both the passing and positioning skills can be understood as forms of multi-agent state synchronisation: determining who holds the ball and how agents align relative to one another along each spatial dimension. This intuition is consistent with insights into position alignment in animal collective movement (animal_study) and emergent coordination in human psychology (emergen_coordination). The key contributions of our paper are:
@@
    Adapting the MacDec-POMDP framework (amato2019masmdp) to support joint-option execution;
We illustrate and experimentally evaluate the capabilities of our approach in two benchmark multi-agent collaboration domains: Level-Based Foraging (papoudakis2021lbf) and Overcooked (ruhdorfer2024overcooked). Beyond assessing the benefits of relative options over standard, non-option-enhanced baselines, we test the hypothesis that the proposed state abstraction facilitates the discovery of a more diverse and generalisable set of coordination behaviours, better suited to support teams of agents in downstream tasks compared with other multi-agent option discovery methods.
@@ 2.1 Decentralised Partially-Observable Markov Decision Processes
A Dec-POMDP (bernstein2009decpomdp) is defined as a tuple $\langle{\mathcal{I}},{\mathcal{S}},\{{\mathcal{A}}^{i}\},{\mathcal{T}},{\mathcal{R}},\{\Omega^{i}\},{\mathcal{O}},\gamma\rangle$, where ${\mathcal{I}}=\{1,\ldots,N\}$ is a finite set of agents’ indices, $\mathcal{S}$} the state space and ${\mathcal{A}}={\mathcal{A}}^{1}\times\cdots\times{\mathcal{A}}^{N}$ the joint action space. At every time-step $t$, agent $i$ receives its local observation $o^{i}_{t}\in\Omega^{i}$, where $\Omega=\Omega^{1}\times\cdots\times\Omega^{N}$ is the joint set of observations that was generated according to the observation function ${\mathcal{O}}:{\mathcal{S}}\times{\mathcal{A}}\times\Omega\rightarrow[0,1]$. Agent $i$ then selects an action $a^{i}_{t}\in{\mathcal{A}}^{i}$ according to a policy $\pi^{i}(a^{i}_{t}|h^{i}_{t})$, that is conditioned on the history of its local observations and actions $h^{i}_{t}=(o^{i}_{1},a^{i}_{1},\ldots,o^{i}_{t-1},a^{i}_{t-1},o^{i}_{t})$. Given the joint set of actions at $t$, $a_{t}=\{a^{1}_{t},\ldots,a^{N}_{t}\}$, the environment transitions to a new state $s_{t+1}\in{\mathcal{S}}$ according to the state transition function ${\mathcal{T}}:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\rightarrow[0,1]$ and induces a global reward received by all agents $r_{t}={\mathcal{R}}(s,a)$, where ${\mathcal{R}}:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\mathbb{R}$. The goal is to learn a joint policy $\pi=(\pi^{1},\ldots,\pi^{N})$ that maximises the expected cumulative discounted return $G=\sum_{t=1}^{T}\gamma^{t-1}r_{t}$.
amato2019masmdp extended Dec-POMDPs by integrating single-agent macro-actions (options) within an asynchronous execution scheme. They defined a MacDec-POMDP as the tuple $\langle{\mathcal{I}},{\mathcal{S}},{\mathcal{A}},\{{\mathcal{M}}^{i}\},{\mathcal{T}},{\mathcal{R}},\{{\mathcal{Z}}^{i}\},\{\Omega^{i}\},\{\zeta^{i}\},O\rangle$, where $\langle{\mathcal{I}},{\mathcal{S}},{\mathcal{A}},{\mathcal{T}},{\mathcal{R}},\{\Omega^{i}\},O\rangle$ are the same as in a Dec-POMDP. The primitive action set of each agent $i\in{\mathcal{I}}$, ${\mathcal{A}}^{i}$, is replaced with a finite set of macro-actions ${\mathcal{M}}^{i}$, and ${\mathcal{M}}={\mathcal{M}}^{1}\times\cdots\times{\mathcal{M}}^{N}$ is the joint macro-action set. They also introduced a joint set of macro-observation $\zeta=\zeta^{1}\times\cdots\times\zeta^{N}$, with ${\mathcal{Z}}^{i}:{\mathcal{M}}^{i}\times{\mathcal{S}}\times\zeta^{i}\rightarrow[0,1]$ specifying the macro-observation probability function for each agent $i$. To formalise macro-actions, separate histories are maintained for the two execution levels: $H^{i}_{A}$ is the primitive action–observation history, and $H^{i}_{M}$ is the macro-action–macro-observation history. The joint primitive history is $H_{A}=(H^{1}_{A},\ldots,H^{N}_{A})$, and the joint macro history is $H_{M}=(H^{1}_{M},\ldots,H^{N}_{M})$.
@@ 2.2 Options
sutton1999optionsdef define an option as a tuple $w=\langle I_{w},\pi_{w},\beta_{w}\rangle$, where $I_{w}\subseteq S$ is the initiation set, $\pi_{w}$ is the option policy, and $\beta_{w}:S\rightarrow[0,1]$ is the termination condition. If these components depend only on the current state, the option is referred to as a Markov option. This notion is generalised to semi-Markov options, in which the policy and termination condition may depend on additional information (e.g., state–action–reward histories), or termination may be triggered by external factors such as a fixed $k$-step horizon.
Eigen-option discovery (machado2017optiondisc) estimates the eigenvectors of the combinatorial graph Laplacian corresponding to a state–transition graph, typically via random walks. The eigenvectors of the graph Laplacian, L=DAL=D-A (where D and A are the degree and adjacency matrices, respectively), captures long-term temporal relationships between states and the overall geometry of an MDP (mahadevan2007protovals). Given an eigenvector $e$, the intrinsic reward function for transitioning from state $s$ to $s^{\prime}$ can be computed as $r_{e}(s,s^{\prime})=e_{i}[s^{\prime}]-e_{i}[s]$. Subsequent work (wu2018optiondisc; jinnai2020optiondisc; wang2021optiondisc) extends this framework to non-tabular domains by approximating the eigenvectors through neural networks trained to minimise objectives derived from graph-drawing theory (koren2005graphdrawing). The ALLO method introduced by gomez2024optiondisc further improves robustness to hyperparameters and eigenvector rotations.
@@
n-metric. Given a set $X$ and an integer $n\geq 2$, an n-(hemi)metric (deza2000nmetrics; deza2009nmetrics) is a function $d:X^{n}\rightarrow\mathbb{R}$, that respects the following conditions:
@@
This definition uses $n$, to replace $m+1$ in the original definition of an m-hemimetric (deza2009nmetrics) and $d(x_{1},\ldots,x_{n})^{z}_{i}$ to represent functions on $n$ elements, where element $i$ is replaced $z$.
n-distance. n-distances (mart2011nmetrics; kiss2018ndistance) relax (M3) by setting $d(x_{1},\ldots,x_{n})=0$ only when $x_{1}=\ldots=x_{n}$, providing a direct way of comparing the dissimilarity or separateness for sets with more than two elements.
@@
The elements of $F_{x}$ are then named the Fermat points for the respective list. Based on these definitions, kiss2018ndistance introduce Fermat n-distances as functions $d_{F}:X^{n}\rightarrow\mathbb{R}$ of the form:
@@ 3 Multi-Agent Option-Discovery
We adopt the successor distances method (myers2024tempdistance) for approximating temporal distances and denote the parameterised state distance as $d_{\theta}:{\mathcal{S}}^{*}\times{\mathcal{S}}^{*}\rightarrow\mathbb{R}$. The Fermat encoder $\phi$ and the distance approximator $d_{\theta}$ are trained concomitantly by enforcing a stop-gradient operator on the distance estimator’s parameters ($\theta$) when integrating it in the Fermat encoder objective (Equation 4).
@@ 3.2 Multi-agent option discovery on relative states
We adopt the approach of dunion2023cmid and introduce a discriminator network $D_{\psi}$ trained to directly detect the information excess. Specifically, the discriminator is trained to distinguishing between real triplets $\{s^{i,j}_{-f,t};s^{i,j}_{f,t};z^{i,j}_{f,t}\}$ obtained by pairing factorised single-agent states from a history of joint states $\tau\triangleq(s_{0},\ldots,s_{T})\sim\rho_{\pi}$, and fake triplets obtained by permuting $s^{i,j}_{-f,t}$ while keeping $\{s^{i,j}_{f,t};z^{i,j}_{f,t}\}$ fixed. As in dunion2023cmid, we use the discriminator’s predictions to define a disentanglement penalty when training $d^{F}_{\theta}$. We defer a description of the CMI minimization algorithm to Appendix A.3. A visual comparison of scalar and the multi-dimensional $n$-distances is in Appendix A.5.
We follow the representation-driven option discovery (ROD) cycle from machado2023optiondisc  [^2] We use the ALLO method from gomez2024optiondisc to approximate the eigenvectors of the graph Laplacian and follow the original eigenoptions approach from machado2017optiondisc, where we use a single ROD cycle to generate the whole set of eigenvectors. A motivation for these design choices is in Appendix A.9. to generate the desired set of options, but precede the eigenvector estimation by first representing the joint states as their disentangled multi-dimensional $n$-distance representation. Both the $n$-distance encoder training and joint option discovery are done offline, prior to the generic training. By converting raw joint states into relative representations, we produce options that reflect a range of complex multi-agent alignment behaviours, enabling agents to synchronise along various subsets of their state features. Figure 2 (right) illustrates the resulting eigenvectors in a grid-world setting with coordinate-based states; Appendix A.6 provides a more detailed analysis. For ease of understanding, we fix the positions of N1N-1 agents and visualise the eigenvector from the remaining agent’s perspective. We emphasise that these relative eigenvectors are highly responsive to changes in the joint state due to the re-centering effect around the Fermat state, a property that is much less evident in the eigenvectors discovered on raw joint-states. In this simple domain, the first eigenvector aligns agents along one coordinate axis, while its negation (also a valid eigenvector) aligns them along the other. The subsequent eigenvector promotes alignment across both axes simultaneously, followed by eigenvectors that capture more complex state synchronisation patterns.
@@ 3.3 Adding Joint Options to Dec-POMPDs
We adapt the MacDec-POMDP framework (amato2019masmdp) described in Section 2.1 to support multi-agent macro-actions (joint options). To ensure the correct selection, execution, and termination of joint options in decentralised settings, we impose the following two modeling assumptions:
@@
Inspired by the work on local options (amato2019masmdp), we define joint options through a hierarchical view on agent histories. A joint option is then a tuple ${\mathcal{W}}=\langle I_{\mathcal{W}},\pi_{\mathcal{W}},\beta_{\mathcal{W}},{\mathcal{P}}^{{\mathcal{I}}}_{\mathcal{W}}\rangle$, where the initiation set $I_{\mathcal{W}}\subseteq H^{M}$ and the termination condition $\beta_{\mathcal{W}}:H^{M}\to[0,1]$ are specified in terms of joint macro histories, and the joint option policy $\pi_{\mathcal{W}}=(\pi_{\mathcal{W}}^{1},\ldots,\pi_{\mathcal{W}}^{N})$ is a map from $H_{A}$ to primitive actions. The multi-agent nature of these options requires additionally specifying the subsets of agents for which an option is defined, ${\mathcal{P}}^{{\mathcal{I}}}_{\mathcal{W}}$, where different join-options might include distinct subsets of agents. For homogeneous teams, ${\mathcal{P}}^{{\mathcal{I}}}_{\mathcal{W}}$ can solely be specified by the number of agents required to initiate the option, i.e., ${\mathcal{P}}^{{\mathcal{I}}}_{\mathcal{W}}=\{J\subseteq{\mathcal{I}}\mid|J|=n_{\mathcal{W}}\}$. This formulation generalises local options, which correspond to the special case where $n_{\mathcal{W}}=1$.
To adapt the MacDec-POMDP framework to support the integration of joint options, we first redefine the macro-actions set $M_{i}$ in terms of joint options, and restrict ourselves to full team consensus, $n_{\mathcal{W}}=N$, while treating primitive actions as joint options with $n_{\mathcal{W}}=1$ and immediate termination. When agent $i$ selects an option $M_{i}$ at time step $t$, this counts as a votevote toward the threshold for initiation, $n_{\mathcal{W}}$; if this value is reached, the option is executed and control is transferred to the option policy until termination, otherwise control is returned at $t+1$. Next, we redefine macro-observations $\zeta_{i}$ to incorporate information shared by teammates. Dec-POMDP-Com (amato2016decpomdpbook) extends the standard framework with an explicit communication protocol. Since communication is not our focus, we mimic this step by allowing agents to share their observations directly and defer a more general treatment of communication to future work.

### v2

@@ 1 Introduction
In single-agent reinforcement learning (RL), the options framework (Sutton et al., 1999) is a widely used mechanism for formulating temporally extended actions. That is, options can act as shortcuts between distant regions of the state space during exploration (McGovern and Barto, 2001). Yet, as noted in prior work (Jong et al., 2008), the effectiveness of options is sensitive to many factors, and poorly designed and excessively large sets of options can hinder learning. This makes option discovery, i.e., the automated design of useful options, a challenging problem. Methods based on the eigen-decomposition of the graph Laplacian, such as Eigenoptions (Machado et al., 2017a), have gained traction due to their task-agnostic discovery of options and exploration guarantees (Jinnai et al., 2019b). However, the reliance on the eigenvectors of the state-transition graph Laplacian leads to an excessive number of options being discovered (twice the state count), a problem that is more pronounced in multi-agent systems because the joint state space grows exponentially with the number of agents. Moreover, current Laplacian eigenvector approximators are most effective in estimating small numbers of eigenvectors (Wang et al., 2021; Gomez et al., 2024), risking not identifying many useful options, particularly those that facilitate exploration at various timescales.
We address option discovery for multi-agent systems through a novel inter-agent relative state abstraction. This new state abstraction compresses the joint state space of a group of agents into a compact latent representation centred around the state of maximal alignment among agents we call the Fermat state. Through this abstraction of the joint state, we drastically reduce the number of discovered options while also focusing the discovery process on inter-agent relational dynamics. We then empirically show how this transformation encourages the emergence of highly coordinated behaviours. Our approach builds on the intuition that in the absence of an explicit objective, synchronisation over state features represents a natural basis for coordination. Returning to the ball game example, both the passing and positioning skills can be understood as forms of multi-agent state synchronisation: determining who holds the ball and how agents align relative to one another along each spatial dimension. This intuition is consistent with insights into position alignment in animal collective movement (Herbert-Read, 2016) and emergent coordination in human psychology (Knoblich et al., 2011). The key contributions of our paper are:
@@
    Adapting the MacDec-POMDP framework (Amato et al., 2019) to support joint-option execution;
We illustrate and experimentally evaluate the capabilities of our approach in two benchmark multi-agent collaboration domains: Level-Based Foraging (Papoudakis et al., 2021) and Overcooked (Ruhdorfer et al., 2024). Beyond assessing the benefits of relative options over standard, non-option-enhanced baselines, we test the hypothesis that the proposed state abstraction facilitates the discovery of a more diverse and generalisable set of coordination behaviours, better suited to support teams of agents in downstream tasks compared with other multi-agent option discovery methods.
@@ 2.1 Decentralised Partially-Observable Markov Decision Processes
A Dec-POMDP (Bernstein et al., 2009) is defined as a tuple $\langle{\mathcal{I}},{\mathcal{S}},\{{\mathcal{A}}^{i}\},{\mathcal{T}},{\mathcal{R}},\{\Omega^{i}\},{\mathcal{O}},\gamma\rangle$, where ${\mathcal{I}}=\{1,\ldots,N\}$ is a finite set of agents’ indices, $\mathcal{S}$} the state space and ${\mathcal{A}}={\mathcal{A}}^{1}\times\cdots\times{\mathcal{A}}^{N}$ the joint action space. At every time-step $t$, agent $i$ receives its local observation $o^{i}_{t}\in\Omega^{i}$, where $\Omega=\Omega^{1}\times\cdots\times\Omega^{N}$ is the joint set of observations that was generated according to the observation function ${\mathcal{O}}:{\mathcal{S}}\times{\mathcal{A}}\times\Omega\rightarrow[0,1]$. Agent $i$ then selects an action $a^{i}_{t}\in{\mathcal{A}}^{i}$ according to a policy $\pi^{i}(a^{i}_{t}|h^{i}_{t})$, that is conditioned on the history of its local observations and actions $h^{i}_{t}=(o^{i}_{1},a^{i}_{1},\ldots,o^{i}_{t-1},a^{i}_{t-1},o^{i}_{t})$. Given the joint set of actions at $t$, $a_{t}=\{a^{1}_{t},\ldots,a^{N}_{t}\}$, the environment transitions to a new state $s_{t+1}\in{\mathcal{S}}$ according to the state transition function ${\mathcal{T}}:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\rightarrow[0,1]$ and induces a global reward received by all agents $r_{t}={\mathcal{R}}(s,a)$, where ${\mathcal{R}}:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\mathbb{R}$. The goal is to learn a joint policy $\pi=(\pi^{1},\ldots,\pi^{N})$ that maximises the expected cumulative discounted return $G=\sum_{t=1}^{T}\gamma^{t-1}r_{t}$.
Amato et al. (2019) extended Dec-POMDPs by integrating single-agent macro-actions (options) within an asynchronous execution scheme. They defined a MacDec-POMDP as the tuple $\langle{\mathcal{I}},{\mathcal{S}},{\mathcal{A}},\{{\mathcal{M}}^{i}\},{\mathcal{T}},{\mathcal{R}},\{{\mathcal{Z}}^{i}\},\{\Omega^{i}\},\{\zeta^{i}\},O\rangle$, where $\langle{\mathcal{I}},{\mathcal{S}},{\mathcal{A}},{\mathcal{T}},{\mathcal{R}},\{\Omega^{i}\},O\rangle$ are the same as in a Dec-POMDP. The primitive action set of each agent $i\in{\mathcal{I}}$, ${\mathcal{A}}^{i}$, is replaced with a finite set of macro-actions ${\mathcal{M}}^{i}$, and ${\mathcal{M}}={\mathcal{M}}^{1}\times\cdots\times{\mathcal{M}}^{N}$ is the joint macro-action set. They also introduced a joint set of macro-observation $\zeta=\zeta^{1}\times\cdots\times\zeta^{N}$, with ${\mathcal{Z}}^{i}:{\mathcal{M}}^{i}\times{\mathcal{S}}\times\zeta^{i}\rightarrow[0,1]$ specifying the macro-observation probability function for each agent $i$. To formalise macro-actions, separate histories are maintained for the two execution levels: $H^{i}_{A}$ is the primitive action–observation history, and $H^{i}_{M}$ is the macro-action–macro-observation history. The joint primitive history is $H_{A}=(H^{1}_{A},\ldots,H^{N}_{A})$, and the joint macro history is $H_{M}=(H^{1}_{M},\ldots,H^{N}_{M})$.
@@ 2.2 Options
Sutton et al. (1999) define an option as a tuple $w=\langle I_{w},\pi_{w},\beta_{w}\rangle$, where $I_{w}\subseteq S$ is the initiation set, $\pi_{w}$ is the option policy, and $\beta_{w}:S\rightarrow[0,1]$ is the termination condition. If these components depend only on the current state, the option is referred to as a Markov option. This notion is generalised to semi-Markov options, in which the policy and termination condition may depend on additional information (e.g., state–action–reward histories), or termination may be triggered by external factors such as a fixed $k$-step horizon.
Eigen-option discovery (Machado et al., 2017b) estimates the eigenvectors of the combinatorial graph Laplacian corresponding to a state–transition graph, typically via random walks. The eigenvectors of the graph Laplacian, L=DAL=D-A (where D and A are the degree and adjacency matrices, respectively), captures long-term temporal relationships between states and the overall geometry of an MDP (Mahadevan and Maggioni, 2007). Given an eigenvector $e$, the intrinsic reward function for transitioning from state $s$ to $s^{\prime}$ can be computed as $r_{e}(s,s^{\prime})=e_{i}[s^{\prime}]-e_{i}[s]$. Subsequent work (Wu et al., 2018; Jinnai et al., 2020; Wang et al., 2021) extends this framework to non-tabular domains by approximating the eigenvectors through neural networks trained to minimise objectives derived from graph-drawing theory (Koren, 2005). The ALLO method introduced by Gomez et al. (2024) further improves robustness to hyperparameters and eigenvector rotations.
@@
n-metric. Given a set $X$ and an integer $n\geq 2$, an n-(hemi)metric (Deza and Rosenberg, 2000; Deza and Deza, 2009) is a function $d:X^{n}\rightarrow\mathbb{R}$, that respects the following conditions:
@@
This definition uses $n$, to replace $m+1$ in the original definition of an m-hemimetric (Deza and Deza, 2009) and $d(x_{1},\ldots,x_{n})^{z}_{i}$ to represent functions on $n$ elements, where element $i$ is replaced $z$.
n-distance. n-distances (Martín and Mayor, 2011; Kiss et al., 2018) relax (M3) by setting $d(x_{1},\ldots,x_{n})=0$ only when $x_{1}=\ldots=x_{n}$, providing a direct way of comparing the dissimilarity or separateness for sets with more than two elements.
@@
The elements of $F_{x}$ are then named the Fermat points for the respective list. Based on these definitions, Kiss et al. (2018) introduce Fermat n-distances as functions $d_{F}:X^{n}\rightarrow\mathbb{R}$ of the form:
@@ 3 Multi-Agent Option-Discovery
We adopt the successor distances method (Myers et al., 2024) for approximating temporal distances and denote the parameterised state distance as $d_{\theta}:{\mathcal{S}}^{*}\times{\mathcal{S}}^{*}\rightarrow\mathbb{R}$. The Fermat encoder $\phi$ and the distance approximator $d_{\theta}$ are trained concomitantly by enforcing a stop-gradient operator on the distance estimator’s parameters ($\theta$) when integrating it in the Fermat encoder objective (Equation 4).
@@ 3.2 Multi-agent option discovery on relative states
We adopt the approach of Dunion et al. (2023) and introduce a discriminator network $D_{\psi}$ trained to directly detect the information excess. Specifically, the discriminator is trained to distinguishing between real triplets $\{s^{i,j}_{-f,t};s^{i,j}_{f,t};z^{i,j}_{f,t}\}$ obtained by pairing factorised single-agent states from a history of joint states $\tau\triangleq(s_{0},\ldots,s_{T})\sim\rho_{\pi}$, and fake triplets obtained by permuting $s^{i,j}_{-f,t}$ while keeping $\{s^{i,j}_{f,t};z^{i,j}_{f,t}\}$ fixed. As in Dunion et al. (2023), we use the discriminator’s predictions to define a disentanglement penalty when training $d^{F}_{\theta}$. We defer a description of the CMI minimization algorithm to Appendix A.3. A visual comparison of scalar and the multi-dimensional $n$-distances is in Appendix A.5.
We follow the representation-driven option discovery (ROD) cycle from Machado et al. (2023)  [^2] We use the ALLO method from Gomez et al. (2024) to approximate the eigenvectors of the graph Laplacian and follow the original eigenoptions approach from Machado et al. (2017b), where we use a single ROD cycle to generate the whole set of eigenvectors. A motivation for these design choices is in Appendix A.9. to generate the desired set of options, but precede the eigenvector estimation by first representing the joint states as their disentangled multi-dimensional $n$-distance representation. Both the $n$-distance encoder training and joint option discovery are done offline, prior to the generic training. By converting raw joint states into relative representations, we produce options that reflect a range of complex multi-agent alignment behaviours, enabling agents to synchronise along various subsets of their state features. Figure 2 (right) illustrates the resulting eigenvectors in a grid-world setting with coordinate-based states; Appendix A.6 provides a more detailed analysis. For ease of understanding, we fix the positions of N1N-1 agents and visualise the eigenvector from the remaining agent’s perspective. We emphasise that these relative eigenvectors are highly responsive to changes in the joint state due to the re-centering effect around the Fermat state, a property that is much less evident in the eigenvectors discovered on raw joint-states. In this simple domain, the first eigenvector aligns agents along one coordinate axis, while its negation (also a valid eigenvector) aligns them along the other. The subsequent eigenvector promotes alignment across both axes simultaneously, followed by eigenvectors that capture more complex state synchronisation patterns.
@@ 3.3 Adding Joint Options to Dec-POMPDs
We adapt the MacDec-POMDP framework (Amato et al., 2019) described in Section 2.1 to support multi-agent macro-actions (joint options). To ensure the correct selection, execution, and termination of joint options in decentralised settings, we impose the following two modeling assumptions:
@@
Inspired by the work on local options (Amato et al., 2019), we define joint options through a hierarchical view on agent histories. A joint option is then a tuple ${\mathcal{W}}=\langle I_{\mathcal{W}},\pi_{\mathcal{W}},\beta_{\mathcal{W}},{\mathcal{P}}^{{\mathcal{I}}}_{\mathcal{W}}\rangle$, where the initiation set $I_{\mathcal{W}}\subseteq H^{M}$ and the termination condition $\beta_{\mathcal{W}}:H^{M}\to[0,1]$ are specified in terms of joint macro histories, and the joint option policy $\pi_{\mathcal{W}}=(\pi_{\mathcal{W}}^{1},\ldots,\pi_{\mathcal{W}}^{N})$ is a map from $H_{A}$ to primitive actions. The multi-agent nature of these options requires additionally specifying the subsets of agents for which an option is defined, ${\mathcal{P}}^{{\mathcal{I}}}_{\mathcal{W}}$, where different join-options might include distinct subsets of agents. For homogeneous teams, ${\mathcal{P}}^{{\mathcal{I}}}_{\mathcal{W}}$ can solely be specified by the number of agents required to initiate the option, i.e., ${\mathcal{P}}^{{\mathcal{I}}}_{\mathcal{W}}=\{J\subseteq{\mathcal{I}}\mid|J|=n_{\mathcal{W}}\}$. This formulation generalises local options, which correspond to the special case where $n_{\mathcal{W}}=1$.
To adapt the MacDec-POMDP framework to support the integration of joint options, we first redefine the macro-actions set $M_{i}$ in terms of joint options, and restrict ourselves to full team consensus, $n_{\mathcal{W}}=N$, while treating primitive actions as joint options with $n_{\mathcal{W}}=1$ and immediate termination. When agent $i$ selects an option $M_{i}$ at time step $t$, this counts as a votevote toward the threshold for initiation, $n_{\mathcal{W}}$; if this value is reached, the option is executed and control is transferred to the option policy until termination, otherwise control is returned at $t+1$. Next, we redefine macro-observations $\zeta_{i}$ to incorporate information shared by teammates. Dec-POMDP-Com (Oliehoek and Amato, 2016) extends the standard framework with an explicit communication protocol. Since communication is not our focus, we mimic this step by allowing agents to share their observations directly and defer a more general treatment of communication to future work.